OpenStudy (meilendurcer):

At the beginning of the school year, Jamal could do 32 sit-ups in one minute. At the end of the school year, Jamal could do 48 sit-ups in one minute. Determine the percent of increase in the number of sit-ups Jamal could do from the beginning of the school year to the end of the school year.

OpenStudy (anonymous):

The increase is 48 - 32 = 16, compared to the starting point (by division) 16/32 = .5, which indicates a 50% improvement.

OpenStudy (anonymous):

Another way to consider it would be to say the end condition compared to the beginning condition is 48/32 = 1.5, which is (astonishingly enough) a 50% increase.